Directions: Write a system of equations that can be used to solve the problem. Then solve the system and determine the answer:
The University of Tennesee women's basketball team made 45 field goals in a recent game; some were 2-pointers and some were 3-pointers. How many 2-point baskets were made and how many 3-point baskets were made if Tennessee scored 101 points?

Answer by Paul(988)

Let the 3-pointer game be x
Let the 2-pointer game be y
x+y=45 (solve for y)
y=45-x (subsitution)
Now:2 pointers plus 3pointers = 101 points So,.
2y+3x=101
Subsitute)
2(45-x)+3x=101
90-2x+3x=101
x=101-90
x=11
y=45-11
y=34

Hence, about 34 2-pointers were played and about 11 3-pointers were played.
Paul.